7. A machinist turns on the power on to a grinding wheel at time t 0s. The wheel accelerates uniformly from rest for 10s and reaches theoperating angular speed of 58 rads. The wheel is run at that angular velocity for 30s, and then power is shut off. The wheel slows down uniformly at 1.4 rads2 until the wheel stops. What is the total number of revolutions made by the wheel in this situation ?1510] 8. A machinist turns on the power on to a grinding wheel at time t-o s. The w heel accelerates uniformly from rest for 10s and reaches the operating angular speed of96 rad s. The wheel is run at that angular velocity for 40 s and then power is shut off. The wheel slows down uniformly at 1.5 rads2until the wheel stops. For how long a time after the power is shut oftdoes it take the wheel to stop?l 64 sl
Answer #1


7) wo = 0 , w = 58 rad/s , t = 10 s

from rotationa kinematic equaitons

w = wo + alpha*t

58 = 0 +alpha*10

alpha = 5.8 rad/s^2

w^2-wo^2= 2alpha*theta

58^2 = 2*5.8*theta

theta1 = 290 rad

w = 58 rad/s , t = 30 s

theta = w*t = 58*30

theta2 = 1740 rad

wo = 58 rad/s , alpha = 1.4 rad/s^2

0 - 58^2 = -2*1.4*theta

theta3 = 1201 rad

Theta = theta1+theta2+theta3

= 290+1740+1201 = 3231 rad

theta = 3231/(2*3.14) = 514 rev


8) w = 96 rad/s, wo = 0

alpha = -1.5 rad/s^2

from rotational kinematic equations

w = wo+alpha*t

0 = 96 -1.5*t

t = 64 s